OpenStudy (anonymous):

Colleen and Bill have just purchased a house for $150,000, with the seller holding a second mortgage of $50,000 plus all accrued interest 5 years from now. How much would they pay if 11 1/4% interest was compounded monthly?

OpenStudy (anonymous):

A=P(1+r/n)^rt =(150,000+50,000)(1+.1125/12)(.1125)(5) I dont have a calculator with me:(

OpenStudy (anonymous):

thank you so much! and its okay i have one :)